JBeats produce and sell a product that has variable costs of $33 and a selling price of $68 . Its current sales total $204,000 per month. Fixed manufacturing costs total $25,000 per month and fixed selling and administrative costs total $17,000 per month. The company is considering a proposal that will increase the selling price by 5%, increase the fixed manufacturing costs by 5%, and increase the fixed selling and administrative costs by $3,500.

A. Compute JBeats’s current break-even point in units.

B. Compute JBeats’s margin of safety in dollars.

C. Compute JBeats’ss net income.

D. Compute JBeats’s breakeven point in units assuming they accept the proposal.

E. Compute JBeats’s net income assuming they accept the proposal and sales total 3,300.
Answer #1

Answers

  • A

A

Sale price per unit

$                            68

B

Variable Cost per unit

$                            33

C = A - B

Contribution margin per unit

$                            35

D = 25000 + 17000

Total Fixed Cost

$                    42,000

E = D/C

Break even in units

1200

  • B

A

Current Sales

$                 204,000

B = 1200 x $ 68

Break even sales

$                    81,600

C = A - B

Margin of Safety

$                 122,400

  • C

A

Contribution margin per unit

$                            35

B = 204000 / $ 68

Units sold

3000

C = A x B

Total Contribution margin

$                 105,000

D = 25000 + 17000

Total Fixed Cost

$                    42,000

E = C- D

Net Income

$                    63,000

  • D

A = 68 x 1.05

New Sales price

$                      71.40

B

Variable cost

$                      33.00

C = A - B

New contribution margin per unit

$                      38.40

D = 25000 x 1.05

New fixed manufacturing cost

$                    26,250

E = 17000 + 3500

New Fixed selling expenses

$                    20,500

F = D+E

Total new Fixed Cost

$                    46,750

G = F/C

Break Even point if proposal is accepted

                         1,217

  • E

A

Unit sold

3300

B

New contribution margin per unit

$                      38.40

C = A x B

Total Contribution margin

$                 126,720

D

New total fixed cost

$                    46,750

E = C - D

Net Income

$                    79,970